OpenStudy (anonymous):

Difficult question. Find the value of\[\sqrt{1-\sqrt{\frac{17}{16}-\sqrt{1-\sqrt{\frac{17}{16}-\sqrt{1-\cdots}}}}}\]

OpenStudy (anonymous):

@KingGeorge @UnkleRhaukus @TuringTest @Zarkon @bahrom7893

OpenStudy (anonymous):

I have the solution, by the way. This is like a challenge. :)

hartnn (hartnn):

y^2 = 1-sqrt{17/16-y} (y^2-1)^2 = 17/16-y solve...

OpenStudy (anonymous):

But \(y\) has one definite value as bounded by the equation before taking the square of both sides. So, which one is it?

hartnn (hartnn):

1/2 ?

OpenStudy (anonymous):

That's a good heuristic guess. It's correct. Why?

OpenStudy (unklerhaukus):

\[x=\sqrt{1-\sqrt{\frac{17}{16}-\sqrt{1-\sqrt{\frac{17}{16}-\sqrt{1-\cdots}}}}}\] \[1-x^2={\sqrt{\frac{17}{16}-\sqrt{1-\sqrt{\frac{17}{16}-\sqrt{1-\cdots}}}}}\] \[\left((1-x^2)^2-\frac{17}{16}\right)^2={{\sqrt{1-\sqrt{\frac{17}{16}-\sqrt{1-\cdots}}}}}\] \[\left((1-x^2)^2-\frac{17}{16}\right)^2=x\]

OpenStudy (unklerhaukus):

am i on the right track?

OpenStudy (anonymous):

A polynomial equation that doesn't rigorously determine which of the \(x\)'s are correct. But yes, you and @hartnn are on the right track. From here it's about iteration behavior at certain values.

OpenStudy (unklerhaukus):

i think i accidentally dropped a negative sign

OpenStudy (anonymous):

Yes, you did. Haha, I didn't notice. I just saw the polynomials and went, "k"

OpenStudy (anonymous):

Not that it makes a difference. A mild algebraic rearrangement leads to the same problem.

OpenStudy (unklerhaukus):

\[\left(\frac{17}{16}-(1-x^2)^2\right)^2=x\]\[\left(\frac{1}{16}+2x^2-x^4\right)^2=x\]\[\frac1{16^2}+\frac{x^2}4-\frac{31x^4}{8}-4x^6+x^8=x\]\[\frac1{256}-x+\frac{x^2}4-\frac{31x^4}{8}-4x^6+x^8=0\]

OpenStudy (anonymous):

Ahaha, that's some unnecessary complexity, but ostensibly it seems right. The difficulty is in showing which one of the \(x\)'s it is, as bounded by the one value \(x\) in the original equation.

OpenStudy (anonymous):

cant we just substitute the value in y^2 = 1-sqrt{17/16-y} and see if it is TRUE OR NOT.

OpenStudy (anonymous):

As squaring gives some unnecessary roots.

OpenStudy (unklerhaukus):

\[1-256x+64{x^2}-992x^4-1024x^6+256x^8=0\] \[1-2^8x+2^6{x^2}-31\times2^5x^4-2^{10}x^6+2^8x^8=0\] \[1-2^8x+2^4{(2x)^2}-62\times(2x)^4-2^{4}(2x)^6+(2x)^8=0\]

OpenStudy (anonymous):